OpenStudy (anonymous):

Where does the graph of x2 - x - 2 = y cross the x-axis?

OpenStudy (anonymous):

@aussy123 , please read the code of conduct : http://openstudy.com/code-of-conduct. You should not be giving out answers directly like that

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

0 = x^2 - x - 2 0 = (x + 1)(x - 2) x + 1 = 0 or x - 2 = 0 x = -1 or x = 2 HAPPY!!

myininaya (myininaya):

@aussy123 , @shivam_bhalla was just informing you of the Code of Conduct which we expect all users to follow. Please don't call others nag especially when they are trying to help you. And yes informing someone of the CoC is helping. Thanks.